OpenStudy (anonymous):

show work. Help me please! sin(x-pi/4)=sqrt 2/2(sin x-cos x)

OpenStudy (campbell_st):

use the difference expansion for sin sin(A - B) = sin(A)cos(B) - cos(A)sin(B) you have A = x and B = pi/4

OpenStudy (campbell_st):

you you'll find you have some exact values... cos(pi/4) and sin(pi/4)

OpenStudy (anonymous):

\[R.H.S=\sin x ~*\frac{ \sqrt{2} }{ 2 }-\cos x~*\frac{ \sqrt{2} }{ 2 }=\sin x ~\cos \frac{ \pi }{ 4 }-\cos x~\sin \frac{ \pi }{ 4 }\] =sin ?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you are welcome.